how many soulutions are in this equation
12x + 6 = 5x

Answer by nerdybill(7384)

how many soulutions are in this equation
12x + 6 = 5x
.
Only one:
12x + 6 = 5x
7x + 6 = 0
7x = -6
x = -6/7